For a standardized normal​ distribution, calculate the probabilities below.

a. ​P(z<1.2​)

b. ​P(z≥0.75​)

c. ​P(−1.23<z<1.45​)
Answer #1

a) P( z< 1.2) = 0.885

{ from normal probability table}

b) P(z>=0.75) = 1- P(z<0.75)

= 1- 0.773

= 0.227

c) P(-1.23<z<1.45)= P(z<1.45) - P(z<-1.23)

= 0.926 - 0.109

= 0.817